OpenStudy (anonymous):

What is the solution of 3|x+5|≤6

OpenStudy (anonymous):

try simplifying first

OpenStudy (anonymous):

i got no solution for the problem

OpenStudy (anonymous):

that is false. so first how can we make it |x+5| <= 2

OpenStudy (anonymous):

i solved that and got -7is less than or equal to x less than or equal to -3

OpenStudy (anonymous):

?

OpenStudy (jamesj):

That's right \[ -7 \leq x \leq -3 \]